I need help on this question please. Thank you!

Find the price a purchaser should be willing to pay for the given bond. Assume that the coupon interest is paid twice a year.​

$12,000 bond with coupon rate 6​% that matures in 4 ​years; current interest rate is 5​%

The purchaser should be willing to pay $
Answer #1

Face value = 12000

Coupon payment = 6%*12000/2 = 360 per half year.

Time = 4*2 = 8 half years

Rate of interest = 2.5% per half year.

Price of the bond = PV of the face value + PV of coupon payment

= 12000(1 + 2.5%)^-8 + 360*(1.025^8 - 1)/(0.025*1.025^8)

= $12,430.21